OpenStudy (anonymous):

if |x| > 1 and |y| < 1/2 , compare x/y and y/x please help!!!

OpenStudy (aravindg):

So we can see that x has more value in magnitude when compared to y. Hence |x/y| is definitely greater. y is smaller so |y/x| is lesser.

OpenStudy (aravindg):

Is that explanation enough.?Else we can go for a detailed one using different signs x and y can take.

OpenStudy (anonymous):

but the answer is cant be determined ??? why?? i cant find

OpenStudy (aravindg):

Oh then lets go for detailed case. You see if x>0 y>0 then x/y is always greater than y/x. But if we x<0 or y<0? Well the case is different.

OpenStudy (anonymous):

ok i got it :))))

OpenStudy (aravindg):

Good :)

OpenStudy (anonymous):

thank u my great friend :***

OpenStudy (aravindg):

In short we dont have an exact answer unless its specifically mentioned x>0 y>0 :) yw friend!!

OpenStudy (aravindg):

Let me know if you need help on any other question!

OpenStudy (anonymous):

Angelo ran 6 miles per hour for 3 miles and 8 miles per hour for 2 miles. What was his average speed, in miles per hour?

OpenStudy (anonymous):

it's a very simple problem and shame on me :))) but cant find the right answer.. i think i don't understand the first part correctly :(((

OpenStudy (aravindg):

Average speed=total distance/total time taken

OpenStudy (anonymous):

yeaa i know that

OpenStudy (aravindg):

You just add total distance he ran and add the time he took.

OpenStudy (anonymous):

Total D= 8+6 Total Time = 6 so 14/6

OpenStudy (aravindg):

You made a mistake there! 8 and 6 are the speeds!

OpenStudy (aravindg):

3 and 2 are the distances :)

OpenStudy (aravindg):

total distance=3+2

OpenStudy (aravindg):

Do u follow that?

OpenStudy (anonymous):

ooooohh just im tooo dumb :(((

OpenStudy (aravindg):

No u r not :/

OpenStudy (anonymous):

thanks again :D

OpenStudy (aravindg):

hey did u get the answer? :O

OpenStudy (anonymous):

6 2/3

OpenStudy (aravindg):

I got 20/3

OpenStudy (anonymous):

yea that's the same :))

OpenStudy (aravindg):

oh wow!! :D

OpenStudy (anonymous):

:D

OpenStudy (aravindg):

Then u call yourself dumb? Actually finding time was the hardest part of the question!!

OpenStudy (aravindg):

Too intelligent friend :D

OpenStudy (anonymous):

:D too smart friend and congratulations you reached too 99 smart score :D

OpenStudy (aravindg):

and became a moderator too :) Thanks!

OpenStudy (anonymous):

but i think you are better than open study can value u ;)

OpenStudy (aravindg):

Hey feel free to message me your doubts via facebook. I can look at them when I get time :)

OpenStudy (aravindg):

Thank you :)

OpenStudy (anonymous):

yea my friend :))) thanks toooo much

OpenStudy (aravindg):

It's my pleasure!

OpenStudy (anonymous):

^_^